5. The probability that a woman does not know swimming is 2/5. If seven women in a city are 

selected at random, find probability that

(i) four women know swimming.

(ii) at least one woman knows swimming


1
Expert's answer
2022-05-23T23:40:55-0400

 Let "X=" the number of female who knows swimming : "X\sim Bin (n, p)."

Given "n=7, q=2\/5=0.4, p=1-q=0.6."

(i)


"P(X=4)=\dbinom{7}{4}(0.6)^4(0.4)^{7-4}=0.290304"



(ii)


"P(X\ge 1)=1-P(X=0)""=1-\dbinom{7}{0}(0.6)^0(0.4)^{7-0}=0.9983616"